Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

mathematics

Expert replies
by Diophantus » Tue Sep 19, 2017 10:51 pm
Three bags labelled P, Q and R contains red, blue and white balls respectively of equal sizes. The ratio of the balls are P:Q=2:3 and Q:R=4:5. All the balls are removed into a big bag and properly mixed together. 1.Find the probability of picking a red ball
2. If two balls are picked at random one after the other with replacement, find the probability of picking :
a. A white ball and a blue ball;
b. A blue ball first and then a red ball.
Join the discussion
Source: — Problem Solving |

by ErikaPrepScholar » Wed Sep 20, 2017 6:59 am
The key thing to notice here is that we are dealing with probability with replacement. Each time we take a ball out of the bag, we put it back in the bag before taking out another ball. This means that the two events are independent — what happens on the first draw doesn't affect the probability of the second draw.

Our first step is to combine our ratios — we know how bag P compares to Q and how Q compares to R, but we don't know how P compares to R. The way we do this is by making Q the same number in each ratio.

Ratios work just like fractions in that we can multiply both sides of the ratio by the same number and still have the same ratio — 1:2 is the same as 2:4 is the same as 3:6 is the same as 150:300.

Our ratios are

P:Q
2:3

Q:R
4:5

The Least Common Multiple of 3 and 4 is 12, so let's make Q = 12 in both ratios.

P:Q
2:3 --- 2*4:3*4 --- 8:12

Q:R
4:5 --- 4:*3:5:*3 --- 12:15

Since Q is now the same in the two ratios, we can combine them.

P:Q:R
8:12:15

Since we don't really care which bags the balls came from, let's rewrite this in terms of color.

red:blue:white
8:12:15

1. The probability of picking a red ball out of all of the balls (all red, blue, and white balls) is 8 / (8+12+15) = 8 / 35.

2a. There are two possibilities for how we can pick a white ball and a blue ball: 1) we can pick a white ball first and then a blue ball, and 2) we can pick a blue ball first and then a white ball. The probability of 1) is 15/35 * 12/35 = 3/7 * 12/35 = 36/245, and the probability of 2) is 12/35 * 15/35 (the reverse) = 36/245. Adding the two probabilities together gives 36/245 + 36/245 = 72/245.

2b. Specifying the order in which the balls are picked limits our possibilities. There is a 12/35 chance we will pick a blue ball first and an 8/35 chance we will pick a red ball second — 12/35 * 8/35 = 96/1225.
Image

Erika John - Content Manager/Lead Instructor
https://gmat.prepscholar.com/gmat/s/

Get tutoring from me or another PrepScholar GMAT expert: https://gmat.prepscholar.com/gmat/s/tutoring/

Learn about our exclusive savings for BTG members (up to 25% off) and our 5 day free trial

Check out our PrepScholar GMAT YouTube channel, and read our expert guides on the PrepScholar GMAT blog
Join the discussion